Two speakers produce identical harmonic waves as i
Answer #1

phase difference \triangle \phi=90

basing on the concept of waves and oscillations

\triangle \phi =2\pi /\lambda (\triangle d)

90 =2\pi /\lambda (\triangle d)

90\lambda/2\pi = (\triangle d)

therefore the mionimum distances (\triangle d)=\lambda /4